A job placement director claims that the average starting salary for nurses is P12,000. A sample of 10 nurses has a mean of P11,750 and a standard deviation of 200. Is there enough evidence to reject the director’s claim at alpha = 0.05

1
Expert's answer
2022-05-26T11:13:49-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=12000"

"H_1:\mu\not=12000"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=9" and the critical value for a two-tailed test is "t_c =2.262156."

The rejection region for this two-tailed test is "R = \{t:|t|>2.262156\}."

The t-statistic is computed as follows:




"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{11750-12000}{200/\sqrt{10}}=-3.9528"

Since it is observed that "|t|=3.9528>2.262156=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=9" degrees of freedom, "t=-3.9528" is "p= 0.003341," and since "p= 0.003341<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 12000, at the "\alpha = 0.05" significance level.